OpenStudy (brittany):

solve using the quadratic formula x^2+5x=6 with steps please

OpenStudy (anonymous):

\[x^2+5x-6=0\] now use quadratic formula

OpenStudy (anonymous):

or you can factor as \[(x+6)(x-1)=0\] set each factor equal 0 and solve

OpenStudy (anonymous):

you get \[x+6=0\] \[x=-6\] or \[x-1=0\] \[x=1\] so your two answers are -6 or 1

OpenStudy (anonymous):

\[x^{2}+5\times x -6=0 \] now by the use of quadratic formula we get \[x=( -b \pm \sqrt{b^2 - 4 \times a \times c})/(2 \times a )\] where a=1, b=5 and c=-6, substiute all these in the above equation we get \[x= (-5 \pm 7)/2\] or, \[x=(-5+7)/2\] and \[x= (-5-7)/2\] on solving this we get x=1 or -6